OpenStudy (anonymous):

roots of the function on these two f(x) = x2 - 4 g(x) = x2 - x - 12

OpenStudy (dumbcow):

f(x) = x^2 -4 = (x+2)(x-2) = 0 x = 2, -2 g(x) = x^2 -x -12 = (x-4)(x+3) = 0 x = 4, -3